OpenStudy (anonymous):

determine whether the points lie on a straight line: a(2,4,2), b(3,7,-2) c(1,3,3) & d(0,-5,5), e (1,-2,4), f(3,4,2) I'm not sure how to solve this, any hints would be appreciated.

OpenStudy (anonymous):

If the slope between every two points is the same, then it is linear

OpenStudy (anonymous):

Slope in XY plane 2-3 / 4-7 = 1/3 Slope in YZ plane 4-7 / 2+2 = -3/4 Slope in XZ plane -1/4

OpenStudy (anonymous):

i see now, thank you!